OpenStudy (anonymous):

In order to prototype a two dimensional game, you first need to create a cardboard mock-up. Since you’ll be creating several versions of this game, you want to save money. Therefore, you have to minimize the amount of cardboard used. The play area of the prototype must be 36 square inches. As shown below, the bottom 1.5 inches of the mock-up is reserved for buttons and controls and the right-hand 2 inches is reserved for information and statistics. What are the total dimensions of the mock-up such that the amount of cardboard used is minimized?

OpenStudy (anonymous):

a. Primary equation Total Area w + 2 * h + 1.5 b. Secondary equation The play area w * h = 36 c. Function of one variable w * h / h = 36 / h w = 36 / h 36 / h + 2 * h + 1.5 d. Feasible domain e. Graph of single variable vs. area f. Final dimensions calculation through differentiation g. Pseudo code of your solution

OpenStudy (anonymous):

I really just need help with the domain to start and I will re-post the others. I would really like help understanding the problems rather than just getting the answer.